EXAMPLE:
A horizontal rope is tied to a 58.0kg box on frictionless ice. What is the tension in the rope if:
1)The box is at rest?
Sol: T=0 N

2)The box moves at a steady vx = 6.00m/s ?
Sol: T=0 N

3)The box vx= 6.00m/s and ax= 5.60m/s^2 ?
Sol: T=325N

QUESTION:
A horizontal rope is tied to a 49.0kg box on frictionless ice. What is the tension in the rope if:
1)The box is at rest?
Sol: T=0 N

2)The box moves at a steady vx= 4.20m/s ?
Sol: T=0 N

3)The box vx= 4.20m/s and ax= 5.60m/s^2 ?
Sol: T=????????????N

$$NOTE:plz answer the question (part3). I already did the first two, so I just need part3. I want to make it easier for you so that I post an example with its answer. you notice its the same question but different numbers, so plz answer the question correctly
Answer #1
Concepts and reason

The concepts used to solve this problem is Newton second of motion.

First, write the expression of force using Newton second of motion then calculate the tension on the rope.

Fundamentals

According to Newton second law of motion, the acceleration of any object which is produced by net force is directly proportional to the magnitude of the force and inversely proportional to the mass of the object.

The expression of tension or force.

T=maT = ma

Here, TTis tension, mm is mass and aa is acceleration.

(3)

Calculate the tension on the rope.

The expression for the tension is,

T=maT = ma

Here, TTis tension, mm is mass and aa is acceleration.

Substitute 49.0kg49.0{\rm{ kg}} for mm and 5.6m/s25.6{\rm{ m/}}{{\rm{s}}^2} for aa.

T=(49.0kg)(5.6m/s2)=274.4N\begin{array}{c}\\T = \left( {49.0{\rm{ kg}}} \right)\left( {5.6{\rm{ m/}}{{\rm{s}}^2}} \right)\\\\ = 274.4{\rm{ N}}\\\end{array}

Ans: Part 3

The tension on the rope is 274N274{\rm{ N}}.